OpenStudy (anonymous):

(_____)^2 = (csc x-1)(csc x+1) What is suppose to be in the blank? ***I know this is a trig identity and do not have my sheet with me. No online refrences are helping me***

OpenStudy (anonymous):

I know it has to be either tan x or cot x. My memory is not on my side today.

OpenStudy (usukidoll):

ok.. there are three trig identities. the most common is \[\cos^2x+\sin^2x=1\]

OpenStudy (usukidoll):

but there are two more equations.

OpenStudy (anonymous):

Ok

OpenStudy (usukidoll):

\[1+\tan^2x=\sec^2x\] \[1+\cot^2x = \csc^2x \]

OpenStudy (usukidoll):

let's solve the right hand side of the equation to see who is the culprit lol xD so we expand \[(cscx-1)(cscx+1)\]

OpenStudy (anonymous):

Haha ok

OpenStudy (usukidoll):

just use foil ... you will notice that O and I cancel out

OpenStudy (anonymous):

Not familiar with FOIL (Sorry). My teacher calls it the "F word" (lol)

OpenStudy (usukidoll):

first outer inner last

OpenStudy (usukidoll):

\[(cscx-1)(cscx+1) = (cscx)(cscx)+(1)(cscx)+(-1)(cscx)+(1)(1)\]

OpenStudy (usukidoll):

fffffffff missed a sign on the last one should be +(-1)(1)

OpenStudy (anonymous):

Oh ok! I think I see what you are getting at

OpenStudy (usukidoll):

\[(cscx-1)(cscx+1) = (cscx)(cscx)+(1)(cscx)+(-1)(cscx)+(-1)(1) = (cscx)^2+cscx-cscx-1\]

OpenStudy (usukidoll):

OY ! \[(cscx)^2+cscx-cscx-1 \]

OpenStudy (usukidoll):

the cscx-cscx cancels out

OpenStudy (usukidoll):

\[\csc^2x-1\]

OpenStudy (anonymous):

Then we can add the 1 right?

OpenStudy (usukidoll):

do you remember \[1+\cot^2x = \csc^2x ? \]

OpenStudy (usukidoll):

what do I need to do have \[\cot^2x \] by itself

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

subract the 1?

OpenStudy (usukidoll):

yes subtract on both sides..

OpenStudy (anonymous):

Oh ok! So we just proved that cot^2 (x) is my answer right?

OpenStudy (usukidoll):

yeah

OpenStudy (usukidoll):

\[\cot^2x = \csc^2x-1\]

OpenStudy (anonymous):

Would I just write it as cotx?

OpenStudy (usukidoll):

huh? you mean for (cotx)^2 = (csc^2x-1)

OpenStudy (anonymous):

Yes

OpenStudy (usukidoll):

hmmm... if you don't forget this part \[\cot^2x = \csc^2x-1 \]

OpenStudy (usukidoll):

\[(cotx)^2 = (cscx)^2-1\] means the same

OpenStudy (usukidoll):

just don't write cot2x and csc2x-1 BIG NO NO!

OpenStudy (anonymous):

Ok! Thank you for your time and teaching me!

OpenStudy (anonymous):

if u observe the RHS, u see it results csc^2-1 there is also an identity saying that, cosec^2-cot^2=1 that implies cosec^2-1=cot^2...therefore theRHS is cot^2 therefore obviously LHS must be cot^2...therefore the answer is cot

OpenStudy (usukidoll):

^ done already lol

OpenStudy (usukidoll):

the term cosecant isn't used for the Pythagorean identities.

OpenStudy (usukidoll):

but there is one identity that is used beyond trig and that's the \[\cos^2x+\sin^2x = 1 \]